OpenStudy (anonymous):

f(x) = (4 − x)^(1/2) what is the inverse if x is less than or equal to 0?

jimthompson5910 (jim_thompson5910):

f(x) = (4 − x)^(1/2) f(x) = sqrt(4 − x) y = sqrt(4 − x) x = sqrt(4 − y) x^2 = 4 - y x^2-4 = -y -x^2+4 = y 4-x^2 = y y = 4-x^2 Inverse is g(x) = 4-x^2 where x >= 0

OpenStudy (anonymous):

would the answer be any different if x </=0 ?

jimthompson5910 (jim_thompson5910):

oh wait, yes it would, one sec

jimthompson5910 (jim_thompson5910):

Inverse is g(x) = 4-x^2 where x >= 2

OpenStudy (anonymous):

so it just changes the way the graph looks?

OpenStudy (anonymous):

the domain

jimthompson5910 (jim_thompson5910):

why x >= 2, because the range of f(x) is y >= 2 Remember that the domain and range swap when it comes to inverses

jimthompson5910 (jim_thompson5910):

yes it changes the way it looks because different pieces are being drawn

OpenStudy (anonymous):

ok i understand, thanks again jim!!